OpenStudy (anonymous):

5m^2 + 32m - 64 over m+8

OpenStudy (anonymous):

First, you need to factor 5m^2 + 32m - 64 Do you want me to explain how it's done?

OpenStudy (anonymous):

you dont factor you have divide m+8 to 5m^2 etc

OpenStudy (anonymous):

5m² + 32 - 63 : m + 8 = 5m - 8 5m² + 40m --------- -8m - 63 -8m -64 ------ 1 Rest Solution: 5m -8 + 1/(m + 8)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

This is polynomial long division, right? This is what I thought the shortcut would be: 5m^2 + 32m - 64 / (m+8) = (5m - 8)(m+8) / (m+8) = 5m - 8 because the (m + 8) cancel. So the solution is 5m - 8. Not sure where the - 63 is coming from. May someone explain?

OpenStudy (anonymous):

yeah i dont know! i think they like to trick people